# AiXpand AI OS - Decentralized ubiquitous computing MLOps execution engine

### v1

@@ 3.7 AiXpand Rapid Application Development
The AiXpand Ecosystem envisions a future where AI-powered application development is accessible to individuals, regardless of their familiarity with programming languages. To achieve this, we are creating tools that enable both developers and non-developers to effortlessly articulate their business requirements and translate them into instructions compatible with the AiXpand SDKs.
@@
For users with existing programming experience, our initial set of tools is designed to facilitate expressing their business needs. We provide code samples that demonstrate the functionality of the AiXpand Ecosystem. To streamline the process, we offer a code-sample gallery tool, allowing users to swiftly run samples, modify code and inputs, and test various results without needing to launch an IDE or create a new project.
@@ 4 Applications
In this section of the paper, we present a range of applications that have been effectively implemented within the AiXpand Network. These applications cover diverse domains and industries, catering to both our internal usage and our customers’ requirements. Additionally, we have collaborated with external partners to develop applications that are specifically tailored for their needs.
While most of these use cases are presently deployed and operational across multiple private or public worker pools that utilize AiXpand E2 on GPU systems, the distribution of workload does not currently incorporate PoAI-based incentivization.
@@ 4.2.2 Stock prediction
An intriguing application of this fusion between AI and everyday operations is the inclusion of prediction functionality within a spreadsheet. To illustrate this application, we can demonstrate how the capabilities of the AiXpand Network can be leveraged. A user selects a specific range of cells within the spreadsheet that contains relevant values for the desired prediction. These values are then transmitted to an API that incorporates AiXpand Nodes. The API performs validation on the received data, formats it for communication with the nearest AiXpand Node, and awaits the response generated by the predictive model. Finally, the response is relayed back to the spreadsheet for display.
The communication between the spreadsheet and the API can be achieved through a RESTful architecture or in real-time using a duplex channel such as websockets. The API utilizes the AiXpand Node.js SDK and taps into the predictive models published within the network to generate the results that are then relayed back to the user of the spreadsheet. For a more detailed depiction of the underlying processes, refer to the swimlane diagram provided in Figure4.
Figure 4: AiXpand network integration in a Google Sheets plugin
@@ 4.3 Video Anonymization - Motionmask™
Motionmask™ illustrates the integration of the AiXpand Network within the backend infrastructure of a traditional Web2 service. The application serves as a web-based tool for anonymizing videos and ensuring compliance with data protection regulations through the provision of an user-friendly interface for identifying and masking specific individuals or vehicles. In order to have a clear view, the swimlane diagram from Figure5 describes the underlying processes in more detail.
Figure 5: AiXpand Network integration in Motionmask™information flows
At the core of Motionmask™ lie very powerful AI detection models. These models require a lot of computing resources which, since Motionmask™ is an online available tool, would have to be provided solely by Motionmask™ itself. This would not be possible and efficient without the AiXpand Network because it would require for Motionmask™ to have an infrastructure small enough to avoid wastage, but also big enough to reliably support surges in user requests.
The interface, which operates within a web browser, communicates with the backend by making RESTful API calls. The backend, in turn, manages all the application business logic such as handling user creation and deletion, registering requests for video anonymization and handling file uploads. The integration of the AiXpand Network is done within a dedicated internal service that connects to the provided MQTT broker and communicates with the nearest AiXpand Node.
Progress is tracked by witnessing the events emitted by the AiXpand Node which describe the state of the running pipeline. Upon completion of processing, the backend services publish an event that is propagated back to the user interface, and the resultant output is presented to the user.
@@ 4.4.1 Decentralized game engine
In the previous examples, the primary focus was on use-cases that employ artificial intelligence particularly shallow and deep machine learning models. Nevertheless, the ecosystem is also capable of supporting the development, integration and even migration of non-AI applications given the hypothesis that the particular target application is suitable for running in a decentralized environment. To demonstrate this versatility, a simple sandbox simulation life-game was designed showcasing how virtual agents interact with one another in an evolutionary heuristic environment. While the initial purpose of this sandbox game was to illustrate the flexibility of AiXpand Development Ecosystem, this mini-project can be used as a tutorial or even as a basic application template.
While being based on the architecture presented in Figure 2, most of the development effort has been geared towards three areas: (i) sandbox simulation encapsulated as a AiXpand E2 model serving module,(ii) the business logic that prepares the pre-rendering vector data and finally (iii) the UI that renders and enables interaction with the sandbox game. The model serving component implements the life-game simulation stateful process itself instead of a traditional stateless machine learning model. While in a conventional model serving process the inferred data is based on input data generated from an upstream data acquisitor, the simulation process here produces the state of the environment after each cycle. The business heuristics component is then responsible for preparing the data that will be delivered from AiXpand E2 to an external custom user interface. Furthermore, the custom UI consumes the inflow of data while generating interaction events in turn that are delivered to the AiXpand Processing Node which runs the simulation game.
@@ 5 Ongoing work
Our development teams are currently working on several crucial areas aimed at addressing key aspects of decentralized execution. These areas include optimizing AI inference and training processes by leveraging homomorphic encryption through our AiXpand EDIL Framework (EDIL - Encrypted Decentralized Inference and Learning) . Our framework currently supports popular deep learning frameworks such as PyTorch and TensorFlow. Additionally, we are actively exploring the utilization of zero-knowledge proofs for evaluating job completion and refining essential workflows involving two-blockchains. This includes establishing efficient interactions between on-chain systems and the AiXpand E2 environment through oracles-based mechanisms. These ongoing developments demonstrate our commitment to advancing decentralized execution capabilities and enhancing the overall functionality and security of the AiXpand Ecosystem.
@@ 5.1 Distributed Homomorphic Encryption
Considering the computational demands of MLOps jobs, it is crucial to minimize the overhead introduced by homomorphic encryption to ensure scalability. To address this concern, our proposed DHE approach utilizes a Domain Auto-encoder-based method, which reduces the amount of information needed while retaining essential data features. From the perspective of a AiXpand Worker, the encoder functions as a one-way function without the decoder, preserving the confidentiality of the data. However, to maintain minimal overhead and achieve the desired secrecy, a unique auto-encoder is required for each job type. This approach allows for secure and efficient manipulation of sensitive data within a decentralized ML environment.
@@ 5.1.1 Decentralized Federated Learning
To dispatch a training job to AiXpand Worker Nodes, the self-designated AiXpand Main Initiator Node either automatically trains a DHE domain auto-encoder with the data to be processed or uses a predefined one from an existing library within the AiXpand Ecosystem. This generates an encoder compressing the data’s salient features. The Initiator Node can then define the end-to-end training job, including the machine deep learning model that consumes the domain encoder output and sends the compressed and DHE data to each worker-peer. Subsequently, workers utilize this DHE approach to train the model using the encoded information in multiple rounds, similar to FedAvg [48], under the supervision of the AiXpand Initiator Node.
@@ 5.1.2 Encrypted Inference
For deploying an inference job to AiXpand Nodes, the AiXpand Initiator Node or external API first encodes the raw data using a neural model stub or a pre-trained neural encoder in compliance with the proposed DHE approach. The Initiator Node then sends both the Main neural model and the encoded data to the partner nodes accepting the job. The receiving nodes forward the input data through the distributed Main model and execute specific post-processing heuristics and code required by the Initiator Node before sending the final result. Throughout this process, receiver nodes only have access to the encoded format received from the Initiator Node as proposed by DHE. By employing this approach, computational load can be distributed across a peer-to-peer decentralized ecosystem.
@@ 5.2 Two layers of BC
The AiXpand Network utilizes a dual blockchain system, consisting of a public and private blockchain, to derive optimal benefits from micro-transactions in the financial domain as well as implement distributed data transactions immutability, respectively. It is imperative to note that while a single blockchain can potentially achieve these objectives, this approach is not favored as it would result in an undesirable gas cost associated with the commitment protocol for data-only transactions that should accommodate an arbitrary load with no correlated financial burden.
The private blockchain is designed as a proprietary, lightweight system to guarantee operational freedom and minimize its architectural footprint. This blockchain can be employed in isolated, private settings and serves the primary function of enabling synchronization among the participating nodes in the network. This synchronization is crucial for the execution of delegated tasks and the signing of system specifications, both of which are integral to the operation of the AiXpand distributed jobs. These processes enable trust-less operations in a heterogeneous operating environment as we aim to achieve a fully trust-less ecosystem for job execution and rewarding.
@@
5.3 Oracles between on-chain and AiXpand E2 worlds
To attain the required level of automation within the ecosystem for multiple aspects from micro-transactions to node reliability history, the deployment of multiple oracles is necessary for the facilitation of communication between the on-chain and AiXpand E2. The oracles enable off-chain to on-chain data transfer between the private and public blockchains in order to enable the allocation of jobs toAixpand Worker Nodes and the creation of worker incentivisation escrow pools. Furthermore, they enable sending commands to worker nodes and retrieve results, assist in computing proofs of job execution and enable on-chain publishing of the outcomes. As central components of the automated end-to-end process, oracles must be secured to prevent the occurrence of vulnerabilities that may have severe implications for the system. It is therefore of utmost importance to ensure that oracles behave in a fair and correct manner, even in scenarios where attackers attempt to compromise the system.
@@ 5.4 Zero Knowledge Proof
In order to enable trust-less data consumption and jobs within the decentralized distributed job execution ecosystem, it is crucial to implement a minimal Zero Knowledge Proof [35] framework and capability in the system. The use of zero-knowledge proof allows the verification of the progress made by the AiXpand Worker Nodes without disclosing any sensitive information regarding the progress of the job, where one only needs to verify that it is increasing, as well as allowing the job initiator to verify the actual completeness of a given finalized job in a feasible time. This not only maintains the confidentiality of the processed data but also minimizes network traffic by reducing the payload exchanged between the Worker and Main Nodes. In this way, the proposed system aims to provide users with a secure and efficient network for the execution of various types of distributed jobs.
@@ 5.5 Conclusion
The vision of the AiXpand revolves around democratizing AI-powered application development and deployment as a transformative ecosystem to address multifaceted challenges and facilitate innovative solutions. All of our tools and technologies streamline the process of converting business requirements into instructions, enabling individuals with diverse backgrounds to seamlessly undergo this transformation.
In a broader context, the AiXpand Ecosystem embodies the remarkable potential of blockchain technology in establishing a more efficient, and inclusive ecosystem, extending beyond the realm of AI. Through its decentralized nature, the ecosystem strives to promote greater autonomy, resilience, and equitable participation, offering a promising paradigm for future advancements in various domains.
AiXpand serves as a testament to the potential of converting computing devices to real assets, while being a pioneer of AI democratization.

### v2

@@ 3.7 Naeural Rapid Application Development
The Naeural Ecosystem envisions a future where AI-powered application development is accessible to individuals, regardless of their familiarity with programming languages. To achieve this, we are creating tools that enable both developers and non-developers to effortlessly articulate their business requirements and translate them into instructions compatible with the Naeural SDKs.
@@
For users with existing programming experience, our initial set of tools is designed to facilitate expressing their business needs. We provide code samples that demonstrate the functionality of the Naeural Ecosystem. To streamline the process, we offer a code-sample gallery tool, allowing users to swiftly run samples, modify code and inputs, and test various results without needing to launch an IDE or create a new project.
@@ 4 Applications
In this section of the paper, we present a range of applications that have been effectively implemented within the Naeural Network. These applications cover diverse domains and industries, catering to both our internal usage and our customers’ requirements. Additionally, we have collaborated with external partners to develop applications that are specifically tailored for their needs.
While most of these use cases are presently deployed and operational across multiple private or public worker pools that utilize Naeural E2 on GPU systems, the distribution of workload does not currently incorporate PoAI-based incentivization.
@@ 4.2.2 Stock prediction
An intriguing application of this fusion between AI and everyday operations is the inclusion of prediction functionality within a spreadsheet. To illustrate this application, we can demonstrate how the capabilities of the Naeural Network can be leveraged. A user selects a specific range of cells within the spreadsheet that contains relevant values for the desired prediction. These values are then transmitted to an API that incorporates Naeural Nodes. The API performs validation on the received data, formats it for communication with the nearest Naeural Node, and awaits the response generated by the predictive model. Finally, the response is relayed back to the spreadsheet for display.
The communication between the spreadsheet and the API can be achieved through a RESTful architecture or in real-time using a duplex channel such as websockets. The API utilizes the Naeural Node.js SDK and taps into the predictive models published within the network to generate the results that are then relayed back to the user of the spreadsheet. For a more detailed depiction of the underlying processes, refer to the swimlane diagram provided in Figure4.
Figure 4: Naeural network integration in a Google Sheets plugin
@@ 4.3 Video Anonymization - Motionmask™
Motionmask™ illustrates the integration of the Naeural Network within the backend infrastructure of a traditional Web2 service. The application serves as a web-based tool for anonymizing videos and ensuring compliance with data protection regulations through the provision of an user-friendly interface for identifying and masking specific individuals or vehicles. In order to have a clear view, the swimlane diagram from Figure5 describes the underlying processes in more detail.
Figure 5: Naeural Network integration in Motionmask™information flows
At the core of Motionmask™ lie very powerful AI detection models. These models require a lot of computing resources which, since Motionmask™ is an online available tool, would have to be provided solely by Motionmask™ itself. This would not be possible and efficient without the Naeural Network because it would require for Motionmask™ to have an infrastructure small enough to avoid wastage, but also big enough to reliably support surges in user requests.
The interface, which operates within a web browser, communicates with the backend by making RESTful API calls. The backend, in turn, manages all the application business logic such as handling user creation and deletion, registering requests for video anonymization and handling file uploads. The integration of the Naeural Network is done within a dedicated internal service that connects to the provided MQTT broker and communicates with the nearest Naeural Node.
Progress is tracked by witnessing the events emitted by the Naeural Node which describe the state of the running pipeline. Upon completion of processing, the backend services publish an event that is propagated back to the user interface, and the resultant output is presented to the user.
@@ 4.4.1 Decentralized game engine
In the previous examples, the primary focus was on use-cases that employ artificial intelligence particularly shallow and deep machine learning models. Nevertheless, the ecosystem is also capable of supporting the development, integration and even migration of non-AI applications given the hypothesis that the particular target application is suitable for running in a decentralized environment. To demonstrate this versatility, a simple sandbox simulation life-game was designed showcasing how virtual agents interact with one another in an evolutionary heuristic environment. While the initial purpose of this sandbox game was to illustrate the flexibility of Naeural Development Ecosystem, this mini-project can be used as a tutorial or even as a basic application template.
While being based on the architecture presented in Figure 2, most of the development effort has been geared towards three areas: (i) sandbox simulation encapsulated as a Naeural E2 model serving module,(ii) the business logic that prepares the pre-rendering vector data and finally (iii) the UI that renders and enables interaction with the sandbox game. The model serving component implements the life-game simulation stateful process itself instead of a traditional stateless machine learning model. While in a conventional model serving process the inferred data is based on input data generated from an upstream data acquisitor, the simulation process here produces the state of the environment after each cycle. The business heuristics component is then responsible for preparing the data that will be delivered from Naeural E2 to an external custom user interface. Furthermore, the custom UI consumes the inflow of data while generating interaction events in turn that are delivered to the Naeural Processing Node which runs the simulation game.
@@ 5 Ongoing work
Our development teams are currently working on several crucial areas aimed at addressing key aspects of decentralized execution. These areas include optimizing AI inference and training processes by leveraging homomorphic encryption through our Naeural EDIL Framework (EDIL - Encrypted Decentralized Inference and Learning) . Our framework currently supports popular deep learning frameworks such as PyTorch and TensorFlow. Additionally, we are actively exploring the utilization of zero-knowledge proofs for evaluating job completion and refining essential workflows involving two-blockchains. This includes establishing efficient interactions between on-chain systems and the Naeural E2 environment through oracles-based mechanisms. These ongoing developments demonstrate our commitment to advancing decentralized execution capabilities and enhancing the overall functionality and security of the Naeural Ecosystem.
@@ 5.1 Distributed Homomorphic Encryption
Considering the computational demands of MLOps jobs, it is crucial to minimize the overhead introduced by homomorphic encryption to ensure scalability. To address this concern, our proposed DHE approach utilizes a Domain Auto-encoder-based method, which reduces the amount of information needed while retaining essential data features. From the perspective of a Naeural Worker, the encoder functions as a one-way function without the decoder, preserving the confidentiality of the data. However, to maintain minimal overhead and achieve the desired secrecy, a unique auto-encoder is required for each job type. This approach allows for secure and efficient manipulation of sensitive data within a decentralized ML environment.
@@ 5.1.1 Decentralized Federated Learning
To dispatch a training job to Naeural Worker Nodes, the self-designated Naeural Main Initiator Node either automatically trains a DHE domain auto-encoder with the data to be processed or uses a predefined one from an existing library within the Naeural Ecosystem. This generates an encoder compressing the data’s salient features. The Initiator Node can then define the end-to-end training job, including the machine deep learning model that consumes the domain encoder output and sends the compressed and DHE data to each worker-peer. Subsequently, workers utilize this DHE approach to train the model using the encoded information in multiple rounds, similar to FedAvg [48], under the supervision of the Naeural Initiator Node.
@@ 5.1.2 Encrypted Inference
For deploying an inference job to Naeural Nodes, the Naeural Initiator Node or external API first encodes the raw data using a neural model stub or a pre-trained neural encoder in compliance with the proposed DHE approach. The Initiator Node then sends both the Main neural model and the encoded data to the partner nodes accepting the job. The receiving nodes forward the input data through the distributed Main model and execute specific post-processing heuristics and code required by the Initiator Node before sending the final result. Throughout this process, receiver nodes only have access to the encoded format received from the Initiator Node as proposed by DHE. By employing this approach, computational load can be distributed across a peer-to-peer decentralized ecosystem.
@@ 5.2 Two layers of BC
The Naeural Network utilizes a dual blockchain system, consisting of a public and private blockchain, to derive optimal benefits from micro-transactions in the financial domain as well as implement distributed data transactions immutability, respectively. It is imperative to note that while a single blockchain can potentially achieve these objectives, this approach is not favored as it would result in an undesirable gas cost associated with the commitment protocol for data-only transactions that should accommodate an arbitrary load with no correlated financial burden.
The private blockchain is designed as a proprietary, lightweight system to guarantee operational freedom and minimize its architectural footprint. This blockchain can be employed in isolated, private settings and serves the primary function of enabling synchronization among the participating nodes in the network. This synchronization is crucial for the execution of delegated tasks and the signing of system specifications, both of which are integral to the operation of the Naeural distributed jobs. These processes enable trust-less operations in a heterogeneous operating environment as we aim to achieve a fully trust-less ecosystem for job execution and rewarding.
@@
5.3 Oracles between on-chain and Naeural E2 worlds
To attain the required level of automation within the ecosystem for multiple aspects from micro-transactions to node reliability history, the deployment of multiple oracles is necessary for the facilitation of communication between the on-chain and Naeural E2. The oracles enable off-chain to on-chain data transfer between the private and public blockchains in order to enable the allocation of jobs toNaeural Worker Nodes and the creation of worker incentivisation escrow pools. Furthermore, they enable sending commands to worker nodes and retrieve results, assist in computing proofs of job execution and enable on-chain publishing of the outcomes. As central components of the automated end-to-end process, oracles must be secured to prevent the occurrence of vulnerabilities that may have severe implications for the system. It is therefore of utmost importance to ensure that oracles behave in a fair and correct manner, even in scenarios where attackers attempt to compromise the system.
@@ 5.4 Zero Knowledge Proof
In order to enable trust-less data consumption and jobs within the decentralized distributed job execution ecosystem, it is crucial to implement a minimal Zero Knowledge Proof [35] framework and capability in the system. The use of zero-knowledge proof allows the verification of the progress made by the Naeural Worker Nodes without disclosing any sensitive information regarding the progress of the job, where one only needs to verify that it is increasing, as well as allowing the job initiator to verify the actual completeness of a given finalized job in a feasible time. This not only maintains the confidentiality of the processed data but also minimizes network traffic by reducing the payload exchanged between the Worker and Main Nodes. In this way, the proposed system aims to provide users with a secure and efficient network for the execution of various types of distributed jobs.
@@ 5.5 Conclusion
The vision of the Naeural revolves around democratizing AI-powered application development and deployment as a transformative ecosystem to address multifaceted challenges and facilitate innovative solutions. All of our tools and technologies streamline the process of converting business requirements into instructions, enabling individuals with diverse backgrounds to seamlessly undergo this transformation.
In a broader context, the Naeural Ecosystem embodies the remarkable potential of blockchain technology in establishing a more efficient, and inclusive ecosystem, extending beyond the realm of AI. Through its decentralized nature, the ecosystem strives to promote greater autonomy, resilience, and equitable participation, offering a promising paradigm for future advancements in various domains.
Naeural serves as a testament to the potential of converting computing devices to real assets, while being a pioneer of AI democratization.